OpenStudy (anonymous):

A plane traveled for 5 hours witha tail wind of 12 mph. The return trip against the same wind took 10 hours. Find the speed of the plane in still air.

OpenStudy (ash2326):

Let the speed of plane be v and speed of wind be u , it's given as 12mph Let the distance it travels be x miles so With the wind it covers x in 5 hours We have distance= speed \(\times\) time with the wind speed of plane is (v+12) \[x= (v+12) \times 5\] against the wind it takes 10 hours the speed of plane = v-12 so \[x=(v-12) \times 10\] Divide the two equations we get \[ 1=\frac{(v+12)\times 5}{(v-12)\times 10}\] we get now \[10v-120=5v-60\] \[5v=60\] Speed of plane in still air \[v= 24 mph\]

OpenStudy (ash2326):

sorry made mistake at last v= 12 mph

OpenStudy (ash2326):

hello8 did you get it?

OpenStudy (anonymous):

yes sorry for late reply

OpenStudy (anonymous):

(I don't know why my PC rebels on me now) Anyway: v1 = v + 12, v2 = v -12 5( v +12 ) = 10 (v- 12) 5v = 120 - 60 -> v = 60/5 = 12 km/h